Which of the following pairs of lines are parallel?

A.y=2x-4y and y=3x+5

B.6x+3y=1 and 4x+2y=1

C.x=3 and y=-4

D.3x-2y=5 and 2x+3y= 4


1
Expert's answer
2021-08-11T11:42:59-0400

Answer:

B.6x+3y=1 and 4x+2y=1


Explanation:

Two lines are parallel if their slopes have the same value. 

The functions: f(x)=m1x+b1 and g(x)=m2x+b2 are parallel if m1=m2.


Let's find the slope for the first function:

3y=1-6x

y=(-6x+1)/3

y=-2x+1/3

The slope is m1=-2.   


Let's find the slope for the second function:

4x+2y=1

2y=-4x+1

y=-2x+1/2

The slope is m2=-2.


So, the functions 6x+3y=1 and 4x+2y=1 are parallel.